ThatCreepyBoy:

What is the solution of log2x − 3125 = 3? x = 1 over 3 x = 1 <<<< x = 7 over 3 x = 4

Vocaloid:

log_2_(x) - 3125 = 3?

Vocaloid:

can you take a screenshot b/c the format is unclear

ThatCreepyBoy:

Sure thing ^-^ sorries

ThatCreepyBoy:

https://prntscr.com/g65aiu

Vocaloid:

we can rewrite this as log_(2x-3)_(5^3) = 3 so 2x - 3 = 5 x = ?

ThatCreepyBoy:

4?

Vocaloid:

good so x = 4 = your answer

ThatCreepyBoy:

Okie :D